MDM Observatory 2.4m Telescope Observing Report for 2026 Jan 16 Observer(s): anderson Institution(s): MDM Instrument: Osmos Worked for 10 hours (all night) Conditions were Photometric most of the night. Problems were encountered, see the separate trouble report for details. Seeing and Weather: Beg: Seeing 1.3 arcsec, Clear, Winds Calm Mid: Seeing 1.3 arcsec, Clear, Winds Calm End: Seeing 1.3 arcsec, Clear, Winds Calm Observing Summary: One of the most frustrating nights in a long time. The 1.3 meter TCS and Xmis programs hung and crashed, and I had to do my best to try and help. Lost a few hours just getting the telescope somewhat stowed and the dome closed. It currently is not operational but will work with eric when I can. Hit targets for Rob and Jimin. ------------------------------ Submitted on 2026 Jan 17 [7:00:56]